# STORM: Segment, Track, and Object Re-Localization from a Single 3D Model

### v1

@@ 5 Conclusion
To conclude, we introduced STORM, an automatic annotation-free framework for 6D object pose estimation and tracking that only requires a single 3D model. STORM integrates a supervision-free segmentation module and a robust tracking component. It incorporates an attention-based feature fusion mechanism for handling both input and reference images. We reformulated tracking as a classification task for detecting tracking failures. In our experiments, STORM achieves high detection performance across multiple standard benchmarks, significantly outperforming existing unsupervised methods and even surpassing state-of-the-art supervised baselines, with minimal computational overhead. Future work will extend STORM to multi-object and category-level tracking, incorporate depth multi-modal sensing, or relational reasoning (Shindo et al. 2024) scale semantic prompting to support more complex manipulation tasks.

### v2

@@ 5 Conclusion
To conclude, we introduced STORM, an automatic, annotation-free framework for 6D object pose estimation and tracking that operates purely from one or a few reference images, without requiring any pre-defined 3D models. STORM integrates a supervision-free segmentation module with a robust tracking component, and employs an attention-based feature fusion mechanism to jointly process input and reference images. We further reformulate tracking as a classification task for detecting tracking failures.
In our experiments, STORM achieves strong detection and pose estimation performance across multiple standard benchmarks, significantly outperforming prior methods and even surpassing state-of-the-art baselines, while incurring minimal computational overhead. Future work will extend STORM to multi-object and category-level tracking, incorporate multi-modal depth sensing, and explore relational reasoning (Shindo et al. 2024) and large-scale semantic prompting to support more complex manipulation tasks.